OpenStudy (anonymous):

solve log (x)^2 +1=5 in decimal form

jagr2713 (jagr2713):

answer choice

OpenStudy (anonymous):

either 100, 10, 316.266 , or 2

jagr2713 (jagr2713):

i got x=2,−2

OpenStudy (junyang96):

10^5 = x^(2)+1 x^(2)= 10^(5) - 1 x= +- square root of 10^(5) - 1

OpenStudy (anonymous):

so then the answer would be 316.266?

OpenStudy (junyang96):

yes. It should have two answers. +3160266 and -316.266. Assuming that your log is log to base 10. Since you only have one choice i guess 316.266 is the best answer

OpenStudy (anonymous):

thank you!!